# Entanglement-enabled advantage for learning a bosonic random displacement channel

### v1

@@ Abstract
Most learning tasks studied so far are restricted to discrete-variable (DV) systems. It is natural to ask whether entanglement-enabled advantage can also be realized for learning properties of bosonic continuous-variable (CV) systems. This is particularly interesting and important because CV systems are ubiquitous in nature and have many applications in quantum information science, such as quantum sensing [25, 26, 27, 6, 8]. However, generalizing the results in DV systems to CV systems can be difficult because bosonic systems have infinite-dimensional Hilbert spaces, making it challenging to formulate rigorous results concerning the complexity of learning properties of these systems. Recent progress has been achieved in studies of entanglement-enhanced learning of CV-state characteristic functions [28]; however, the lower bounds obtained so far apply to a restricted class of learning strategies rather than to general entanglement-free schemes.
@@
In this work, we rigorously establish an entanglement-enabled advantage in learning a probabilistic mixture of $n$-mode displacement operations, called a bosonic random displacement channel. Specifically, we show that any schemes without ancillary quantum memory require a number of samples exponential in $n$ to learn the characteristic functions of the channel with reasonably good precision and high success probability. On the contrary, we present a simple scheme utilizing entanglement with ancillary quantum memory (i.e., entanglement-assisted) that can complete the same learning task with a sample complexity independent of $n$, given access to two-mode squeezed vacuum (TMSV) states with sufficiently large squeezing parameter and Bell measurements (BM). This establishes an exponential separation between learning with and without entanglement in the bosonic system. The two learning scenarios are illustrated in Fig. 1. Note that our hardness results hold for arbitrarily high-energy input states and arbitrary measurements, while the presented entanglement-assisted scheme only requires a finite-energy TMSV and BM.
Furthermore, we analyze the robustness of this entanglement-enabled advantage under realistic experimental conditions. Specifically, we study the photon-loss effect, the most common noise source in optical platforms. Our results suggest that for squeezing parameters and loss rates achievable in a state-of-the-art bosonic experiment platform, the separation in sample complexity remains significant. Therefore, we anticipate that an experimental demonstration of entanglement-enabled advantage in CV quantum systems can be achieved in the near future.
@@
where $\hat{D}(\alpha)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\otimes_{i=1}^{n}\hat{D}(\alpha_{i})$ and $\hat{D}(\alpha_{i})\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\exp(\alpha_{i}\hat{a}_{i}^{\dagger}-\alpha_{i}^{*}\hat{a}_{i})$ is the displacement operator for the $i$th mode. The random displacement channel can also be equivalently described by the characteristic function of $p(\alpha)$, i.e., its Fourier transform, as (see SM S1 [29] for the derivation)
@@
Here, because of the Fourier relation, $\lambda(\beta)$ with a large $\beta$ contributes to rapidly oscillating $p(\alpha)$. Since the domain of $\beta$ is infinite in principle, we will focus on a restricted finite domain specified later. The goal is to learn the channel by estimating the characteristic function $\lambda(\beta)$. We emphasize that the goal is to characterize the channel, as opposed to identifying a particular displacement that is drawn from the distribution $p(\alpha)$. The value of $\beta$ for which $\lambda(\beta)$ is to be estimated is revealed only after all measurements are completed.
We focus on the separation between two types of learning schemes for the random displacement channel distinguished by whether or not the scheme uses entanglement between the system and an ancilla, as illustrated in Fig. 1. Throughout this work, we define an entanglement-free scheme to be both ancilla-free and concatenation-free, i.e., the output of the channel is measured destructively after each channel use. However, the entanglement-free scheme is allowed to be adaptive; for each channel use, the input to the channel and the measurement performed on the output may depend on measurement outcomes obtained in earlier rounds. This scenario is similar to Refs. [17, 22]. Several recent works have obtained lower bounds on learning DV channels that hold even with concatenation [23, 30, 31], but we will not analyze the consequences of concatenating CV channels in this work for simplicity.
Figure 1: Schemes for learning an $n$-mode random displacement channel $\Lambda$. (a) TMSV+BM, a specific entanglement-assisted (EA) scheme. (b) General entanglement-free (EF) scheme. Here we assume no concatenation is allowed, i.e., each copy of the channel acts on some input state $\rho_{0}$ and is measured destructively by some POVM $\{E\}$. The input state and measurement are allowed to be adaptively chosen depending on previous outcomes. An example of EF scheme is Vacuum+Heterodyne (see the main text).
Schemes.— Now, we present an entanglement-assisted scheme (see Fig. 1) inspired by a similar scheme to which has been proposed in DV Pauli channel estimation in Ref. [23]. Consider an $n$-mode random displacement channel $\Lambda_{B}$ acting on the $n$-mode system $B$. To learn this channel, we prepare $n$ CV Bell states with a finite squeezing parameter $r$, which is a two-mode squeezed vacuum (TMSV) state, and half of the states go through the channel while the other half stays in quantum memory. Finally, we measure the output state by CV Bell measurement (BM), which can be implemented by passing through a 50:50 beam splitter and performing homodyne measurement on output ports along different quadratures [27]. Formally, the BM POVM element labeled by $\{\zeta\in\mathbb{C}^{n}\}$ has the following form: $(I\otimes\hat{D}(\zeta))|\Psi\rangle\langle\Psi|(I\otimes\hat{D}^{\dagger}(\zeta))/\pi^{n}$; here $|\Psi\rangle$ denotes the tensor product of $n$ infinitely squeezed TMSV states, each proportional to $\sum_{k=0}^{\infty}|k\rangle|k\rangle$ when expressed in the Fock basis. To see how to learn a random displacement channel using this TMSV+BM scheme, we invoke the probability of obtaining outcome $\zeta$ from BM (see SM S2 A [29] for the derivation):
@@
$\lambda(\beta)$ $=e^{e^{-2r}|\beta|^{2}}\int d^{2n}\zeta~{}p_{EA}(\zeta)e^{\zeta^{\dagger}\beta-\beta^{\dagger}\zeta}$ (5)
$\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=e^{e^{-2r}|\beta|^{2}}\lambda_{EA}(\beta).$
This expression indicates that, by sampling $N$ measurement outcomes $\{\zeta^{(i)}\}_{i=1}^{N}$ from a TMSV+BM scheme, one can obtain an unbiased estimator $\tilde{\lambda}(\beta)$ of $\lambda(\beta)$ by defining $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{e^{-2r}|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$. Note that the same set of samples can be used to estimate $\lambda(\beta)$ for different values of $\beta$ just by modifying the estimator. Using the Hoeffding’s bound, we prove the following theorem (see SM S2 A for the proof):
@@
Let us compare the TMSV+BM scheme with a particular entanglement-free scheme that uses the vacuum state as input and heterodyne detection (Vacuum+Heterodyne). Here, heterodyne detection is defined as a projection onto the (overcomplete) basis of coherent states, i.e., $|\zeta\rangle\langle\zeta|/\pi^{n}$ with $\zeta\in\mathbb{C}^{n}$. Though it may not be the optimal entanglement-free scheme, this specific scheme helps us understand the limitations of entanglement-free schemes, which we capture more generally in Theorem 2 below.
In this scheme, the probability of obtaining POVM outcome $\zeta$ is (see SM S2 B [29])
@@
In fact, the Vacuum+Heterodyne scheme can be understood as the TMSV+BM scheme with $r=0$. Inverting this relation by Fourier transforming, we may express the channel’s characteristic function in terms of the measurement probability distribution:
@@
which yields another unbiased estimator $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$ given $N$ samples $\{\zeta^{(i)}\}_{i=1}^{N}$. Comparing to (5), we see that the $r$-dependent prefactor is missing from (7). Specifically, if we confine $\beta$ to $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, we obtain upper bounds on the sample complexity for achieving an error $\epsilon$ with success probability $1{-}\delta$ using each scheme:
$N_{EA}$ $=O(e^{2e^{-2r}\kappa n}\epsilon^{-2}\log\delta^{-1}),$ (8)
$N_{VH}$ $=O(e^{2\kappa n}\epsilon^{-2}\log\delta^{-1}).$ (9)
In particular, if we choose the squeezing parameter as $r=\Omega(\log n)$, the sample complexity $N_{EA}=O(\epsilon^{-2}\log\delta^{-1})$ of the entanglement-assisted scheme becomes independent of the number of modes $n$, while our upper bound on sample complexity of the entanglement-free scheme increases exponentially with $n$. Since the accessible squeezing parameter is bounded in practice, though, we will compare the sample complexities of the two schemes when $r$ is an $n$-independent constant below.
To illustrate the difference, we compare TMSV+BM and Vacuum+Heterodyne strategies with an example in Fig. 2. We consider a single-mode channel for ease of visualization, characterized by
$p(\alpha)$ $=\frac{2\sigma^{2}}{\pi}e^{-2\sigma^{2}|\alpha|^{2}}[\cos^{2}(\alpha_{r}\gamma_{i}-\alpha_{i}\gamma_{r})$
$~{}~{}~{}~{}~{}+\sin^{2}(\alpha_{r}\gamma_{r}+\alpha_{i}\gamma_{i})],$ (10)
@@
with $\sigma=0.3$, $\gamma_{r}=1.6$, $\gamma_{i}=0$ ($\gamma\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\gamma_{r}+i\gamma_{i}$), and $r=2$ for the TMSV+BM scheme. The figure, where we present the underlying output probability distributions and their characteristic functions from Eqs. (4),(5),(6), and (7), clearly shows that in the TMSV+BM scheme with a sufficiently large squeezing parameter, the resultant probability distribution and characteristic function are almost identical to the ideal case. However, for the Vacuum+Heterodyne scheme, the vacuum noise distorts the initial probability distribution so significantly that we cannot see the signal clearly, which thus makes it harder to estimate the original characteristic function.
Figure 2: Comparison between (a) the true distribution, (b) TMSV+BM, and (c) Vacuum+Heterodyne strategies. The left panel represents the probability distribution of the true distribution and measurement probability distributions for each scheme. The right panel represents the characteristic function of probability distributions.
Lower bound.— Our upper bound on the sample complexity of the Vacuum+Heterodyne scheme scales exponentially with $n$. Can this scaling be improved using more advanced entanglement-free schemes, such as homodyne or general-dyne detection [32, 27], or by non-Gaussian resources like GKP states [33] or photon-number resolving measurements [34]? Here, using information-theoretic methods, we prove an exponential sample complexity lower bound for any entanglement-free scheme. This highlights the indispensable role of entanglement for efficiently learning bosonic random displacement channels. Our result is as follows:
@@
Let $\Lambda$ be an arbitrary $n$-mode random displacement channel ($n\geq 8$) and consider an entanglement-free scheme that uses $N$ copies of $\Lambda$. After all measurements are completed, the scheme receives the query $\beta\in\mathbb{C}^{n}$ and returns an estimate $\tilde{\lambda}(\beta)$ of $\Lambda$’s characteristic function $\lambda(\beta)$. Suppose that, with success probability at least $2/3$, $|\tilde{\lambda}(\beta)-\lambda(\beta)|\leq\epsilon\leq 0.24$ for all $\beta$ such that $|\beta|^{2}\leq n\kappa$. Then $N\geq 0.01\epsilon^{-2}(1+1.98\kappa)^{n}$.
Here, the choice of success probability $2/3$ is arbitrary and can be easily amplified. Comparing with the entanglement-assisted sample complexity given in Eq. (8), Theorem 2 establishes a separation exponential in $n$ for cutoff coefficient $\kappa=O(1)$ and squeezing parameter $r=\Omega(\log n)$. The intuition underlying this theorem is that displacement operators $\hat{D}(\beta)$ do not generally commute with each other. Consequently, entanglement-free measurements can resolve $\lambda(\beta)$ for only a small portion of $\beta$ space. We sketch the proof below and leave the full details to SM S3 [29].
@@
Our proof extends the techniques of Refs. [23, 18] to the CV case. We begin by defining the following family of “$3$-peak” random displacement channels $\bm{\Lambda}^{\epsilon,\sigma}_{\mathrm{3\mhyphen peak}}=\{\Lambda_{\gamma}\}_{\gamma\in\mathbb{C}^{n}}$ whose characteristic functions and distributions of displacement are, respectively,
@@
with positive parameters $\sigma$ and $\epsilon\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=0.98\epsilon_{0}$. We will show that, even with the prior knowledge that the channel is from this family, it is still hard for entanglement-free schemes to complete the learning tasks.
The key idea is to reduce learning to binary hypothesis testing. Consider the following game between Alice and Bob: Alice chooses one of two hypotheses with equal probability: (1) Set $\Lambda=\Lambda_{0}$; (2) Set $\Lambda=\Lambda_{\gamma}$, for $\gamma$ sampled from a zero-mean homogeneous Gaussian distribution whose variance is determined by $\kappa$. Next, Alice allows Bob to use the channel $\Lambda$ $N$ times, and Bob uses his favorite entanglement-free scheme to learn from these channel uses. After Bob has finished all quantum measurements and keeps only classical data, Alice reveals some auxiliary information to Bob, who is then asked to decide whether Alice has chosen (1) or (2).
Given a learning scheme satisfying the assumptions of Theorem 2, Bob can guess correctly with high probability. This means that the outcome distributions of Bob’s scheme under hypotheses (1) and (2) must have a sufficiently large total variation distance (TVD). On the other hand, we can upper bound the contribution from each use of $\Lambda$ to the TVD to be exponentially small, where we use a technique inspired by Ref. [35] which derived the maximum fidelity of Gaussian random displacement channels. Therefore, the number of channel uses $N$ must be exponentially large to ensure a large enough TVD, which gives us the desired lower bound. ∎
@@
Effect of loss.— Now, for practical applications, we study how the entanglement-assisted scheme is affected by photon loss, a dominant noise source in optical platforms (see SM S2 D for a discussion of more general noise models, such as phase diffusion). Photon loss transforms the relevant bosonic operator $\hat{a}$ to $\sqrt{T}\hat{a}+\sqrt{1-T}\hat{e}$, where $T$ is the transmission rate and $\hat{e}$ is the environmental mode, i.e., $1-T$ is the loss rate. We consider two different places where the loss occurs: one is before applying the channel with loss rate $1-T_{b}$ to model the preparation imperfection, and the other is after applying the channel and before the perfect BM with loss rate $1-T_{a}$, which models the finite efficiency of detection, i.e., an imperfect BM [27]. As before, we derive the relation between the measurement probability distribution and the characteristic function of the channel (with appropriate rescaling of the phase):
@@
which incorporates the loss rates. Because loss degrades the advantage from squeezing, the upper bound on sample complexity in Theorem 1 is modified in the presence of loss (see SM S2 C for the proof):

### v2

@@ Abstract
Most learning tasks studied so far are restricted to discrete-variable (DV) systems. It is natural to ask whether entanglement-enabled advantage can also be realized for learning properties of bosonic continuous-variable (CV) systems. This is particularly important because CV systems are ubiquitous in nature and have many applications in quantum information science, such as quantum sensing [25, 26, 27, 6, 8]. However, generalizing the results in DV systems to CV systems is challenging because bosonic systems have infinite-dimensional Hilbert spaces, complicating rigorous complexity analysis of learning. Recent progress has been achieved in studies of entanglement-enhanced learning of CV-state characteristic functions [28]; however, the lower bounds obtained so far apply to a restricted class of learning strategies rather than to general entanglement-free schemes. There have also been recent works studying applications of statistical learning theory to CV systems [29, 30, 31].
@@
Characterizing displacements applied to CV systems has many applications, including force sensing [32], dark matter search [33, 34, 35], gravitational wave detection [36], and Raman scattering spectroscopy [37]; therefore this task has drawn substantial recent attention [38, 39]. In this work, we rigorously establish an entanglement-enabled advantage in learning a probabilistic mixture of $n$-mode displacement operations, called a random displacement channel. Specifically, we prove that any schemes without ancillary quantum memory (i.e., entanglement-free) require exponentially many samples in $n$ to learn the channel’s characteristic function with good precision and high success probability. In contrast, we present a simple scheme utilizing entanglement with ancillary quantum memory (i.e., entanglement-assisted) that can achieve the same learning task with sample complexity independent of $n$, given access to two-mode squeezed vacuum (TMSV) states with sufficiently large squeezing parameter and Bell measurements (BM). This establishes an exponential separation between learning with and without entanglement in the bosonic system (see Fig. 1).
Notably, our entanglement-assisted scheme achieves an exponential advantage using only Gaussian input state preparation and measurements which are experimentally feasible. Furthermore, we demonstrate the robustness of the entanglement-enabled advantage under realistic experimental conditions by analyzing the effects of photon loss, phase noise, and crosstalk, the most common noise sources in optical platforms, suggesting that for squeezing parameters and noise rates achievable in a state-of-the-art experiment, the advantage remains significant and thus can be realized shortly.
For DV systems, existing techniques for proving quantum advantages in learning typically involve characterizing the hardness of distinguishing a given state or channel from the maximally mixed state or completely depolarizing channel [17, 18, 19, 23]. These states and channels are ill-defined for CV systems; therefore we needed a different approach to establish an entanglement-enabled advantage without making reference to unphysical states (see the proof of Theorem 2). Another distinctive feature of our results is that our entanglement-free lower bound (Theorem 2) holds even for schemes that use input states with arbitrarily high energy (i.e., occupying an arbitrarily large Hilbert space), while the entanglement-assisted upper bound (Theorem 1) only uses finite energy input states.
@@
where $\hat{D}(\alpha)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\otimes_{i=1}^{n}\hat{D}(\alpha_{i})$ and $\hat{D}(\alpha_{i})\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\exp(\alpha_{i}\hat{a}_{i}^{\dagger}-\alpha_{i}^{*}\hat{a}_{i})$ is the displacement operator for the $i$th mode. It can be equivalently described by the characteristic function of $p(\alpha)$, i.e., its Fourier transform, as (see SM S1 [40] for the derivation)
@@
Here, because of the Fourier relation, $\lambda(\beta)$ with a large $\beta$ contributes to rapidly oscillating $p(\alpha)$. Since the domain of $\beta$ is infinite, we will focus on a restricted finite domain specified later. The goal is to learn the channel by estimating the characteristic function $\lambda(\beta)$; this is distinct from identifying a particular displacement drawn from the distribution $p(\alpha)$. The value of $\beta$ of $\lambda(\beta)$ to be estimated is revealed only after all measurements are completed.
We focus on the separation between two types of learning schemes for the random displacement channel distinguished by whether the scheme uses entanglement between the system and an ancilla (see Fig. 1). Throughout this work, we define an entanglement-free scheme as being both ancilla-free and concatenation-free, i.e., the channel’s output is measured destructively after each channel use. However, it is allowed to be adaptive; for each channel use, the input to the channel and the measurement may depend on measurement outcomes obtained in earlier rounds. This scenario is similar to Refs. [17, 22]. Several recent works have obtained lower bounds on learning DV channels that hold even with concatenation [23, 41, 42], but for simplicity, we focus on concatenation-free schemes.
Figure 1: Schemes for learning an $n$-mode random displacement channel $\Lambda$. (a) TMSV+BM, a specific entanglement-assisted (EA) scheme. (b) General entanglement-free (EF) scheme. See the main text for more details.
Schemes.— Now, we present an entanglement-assisted scheme (see Fig. 1) inspired by Pauli channel learning proposed in Ref. [23]. Consider an $n$-mode random displacement channel $\Lambda_{B}$ acting on the $n$-mode system $B$. To learn this channel, we prepare $n$ CV Bell states with a finite squeezing parameter $r$, i.e. TMSV states, and half of the states go through the channel while the other half stays in quantum memory. Finally, we measure the output state by CV BM, implemented by passing through 50:50 beam splitters and performing homodyne measurement on output ports along different quadratures [27]. Formally, the BM POVM element labeled by $\{\zeta\in\mathbb{C}^{n}\}$ has the following form: $(I\otimes\hat{D}(\zeta))|\Psi\rangle\langle\Psi|(I\otimes\hat{D}^{\dagger}(\zeta))/\pi^{n}$; here $|\Psi\rangle$ denotes the tensor product of $n$ infinitely squeezed TMSV states, each proportional to $\sum_{k=0}^{\infty}|k\rangle|k\rangle$ when expressed in the Fock basis. To see how to learn a random displacement channel using this TMSV+BM scheme, we invoke the probability of obtaining outcome $\zeta$ from BM (see SM S2 A [40] for the derivation):
@@
$\lambda(\beta)$ $=e^{e^{-2r}|\beta|^{2}}\int d^{2n}\zeta~{}p_{EA}(\zeta)e^{\zeta^{\dagger}\beta-\beta^{\dagger}\zeta}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=e^{e^{-2r}|\beta|^{2}}\lambda_{EA}(\beta).$ (5)
This expression indicates that given $N$ outcomes $\{\zeta^{(i)}\}_{i=1}^{N}$ from the TMSV+BM scheme, $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{e^{-2r}|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$ is an unbiased estimator of $\lambda(\beta)$. Note that the same set of samples can be used to estimate $\lambda(\beta)$ for different values of $\beta$ just by modifying the estimator. Using the Hoeffding’s bound, we prove the following theorem (see SM S2 A for the proof):
@@
In Theorem 2 below we formulate a lower bound on the sample complexity for general entanglement-free schemes. First, though, let us compare the entanglement-assisted TMSV+BM scheme with a particular entanglement-free scheme that uses the vacuum state as input and heterodyne detection (Vacuum+Heterodyne). Here, heterodyne detection is defined as a projection onto the (overcomplete) basis of coherent states, i.e., $|\zeta\rangle\langle\zeta|/\pi^{n}$ with $\zeta\in\mathbb{C}^{n}$. This specific scheme helps us to understand the limitations of more general entanglement-free schemes which are captured more fully by Theorem 2.
In this scheme, the probability of obtaining the outcome $\zeta$ is (see SM S2 B [40])
@@
In fact, the Vacuum+Heterodyne scheme can be understood as the TMSV+BM scheme with $r=0$. Inverting this relation by Fourier transforming, we again express the channel’s characteristic function by the measurement probability distribution:
@@
which yields another unbiased estimator $\tilde{\lambda}(\beta)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\frac{1}{N}e^{|\beta|^{2}}\sum_{i=1}^{N}e^{\zeta^{(i)\dagger}\beta-\beta^{\dagger}\zeta^{(i)}}$ given $N$ outcomes $\{\zeta^{(i)}\}_{i=1}^{N}$. Comparing to (5), the $r$-dependent prefactor is missing from (7). Specifically, if we confine $\beta$ to $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, we obtain upper bounds on the sample complexity for achieving an error $\epsilon$ with success probability $1{-}\delta$ using each scheme:
$N_{EA}=O(e^{2e^{-2r}\kappa n}\epsilon^{-2}\log\delta^{-1}),N_{VH}=O(e^{2\kappa n}\epsilon^{-2}\log\delta^{-1}).$ (8)
In particular, if we choose $r=\Omega(\log n)$, the sample complexity of the entanglement-assisted scheme becomes independent of the number of modes $n$, while our upper bound on that of the entanglement-free scheme increases exponentially with $n$. Since the accessible squeezing parameter is bounded in practice, though, we will consider the sample complexities for constant $r$ below.
To illustrate the difference, we compare the two strategies with a single-mode example, characterized by
$p(\alpha)$ $=\frac{2\sigma^{2}}{\pi}e^{-2\sigma^{2}|\alpha|^{2}}[\cos^{2}(\alpha_{r}\gamma_{i}-\alpha_{i}\gamma_{r})+\sin^{2}(\alpha_{r}\gamma_{r}+\alpha_{i}\gamma_{i})],$ (9)
@@
with $\sigma=0.3$, $\gamma_{r}=1.6$, $\gamma_{i}=0$ ($\gamma\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\gamma_{r}+i\gamma_{i}$), and $r=2$ for the TMSV+BM scheme. Figure 2, presenting the underlying output probability distributions and their characteristic functions from Eqs. (4),(5),(6), and (7), clearly shows that in the TMSV+BM scheme with a sufficient squeezing parameter, the resultant probability distribution and characteristic function are almost identical to the ideal case. However, for the Vacuum+Heterodyne scheme, the vacuum noise distorts the initial probability distribution so significantly that we cannot see the signal clearly, making it harder to estimate the original characteristic function.
Figure 2: Comparison between (a) the true distribution, (b) TMSV+BM, and (c) Vacuum+Heterodyne strategies. The left panel represents the probability distribution of the true distribution and measurement probability distributions for each scheme. The right panel represents the characteristic function of the probability distributions.
Lower bound.— Our upper bound on the sample complexity of the Vacuum+Heterodyne scheme scales exponentially with $n$. Can this scaling be improved using more advanced entanglement-free schemes, such as general-dyne detection or photon-number resolving measurements [43], or by non-Gaussian resources like GKP states [44]? Here, using information-theoretic methods, we prove an exponential sample complexity lower bound for any entanglement-free scheme. This highlights the indispensable role of entanglement for efficiently learning random displacement channels. Our result is as follows:
@@
Let $\Lambda$ be an arbitrary $n$-mode random displacement channel ($n\geq 8$) and consider an entanglement-free scheme that uses $N$ copies of $\Lambda$. After all measurements are completed, the scheme receives the query $\beta\in\mathbb{C}^{n}$ and returns an estimate $\tilde{\lambda}(\beta)$ of $\Lambda$’s characteristic function $\lambda(\beta)$. Suppose that, with success probability at least $2/3$, $|\tilde{\lambda}(\beta)-\lambda(\beta)|\leq\epsilon\leq 0.24$ for all $\beta$ such that $|\beta|^{2}\leq n\kappa$ and all $\Lambda$. Then $N\geq 0.01\epsilon^{-2}(1+1.98\kappa)^{n}$.
Here, the choice of success probability $2/3$ is arbitrary and can be easily amplified. Comparing with the entanglement-assisted sample complexity given in Eq. (8), Theorem 2 establishes a separation exponential in $n$ for cutoff coefficient $\kappa=O(1)$ and squeezing parameter $r=\Omega(\log n)$. The intuition underlying this theorem is that displacement operators do not generally commute with each other. Consequently, entanglement-free measurements can resolve $\lambda(\beta)$ for only a small portion of $\beta$ space. We sketch the proof below and leave the full details to SM S3 [40].
@@
We begin by defining the following family of “$3$-peak” random displacement channels $\bm{\Lambda}^{\epsilon,\sigma}_{\mathrm{3\mhyphen peak}}=\{\Lambda_{\gamma}\}_{\gamma\in\mathbb{C}^{n}}$ whose characteristic functions and distributions of displacement are, respectively,
@@
with positive parameters $\sigma$ and $\epsilon\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=0.98\epsilon_{0}$. Note that the channel with $\sigma\to 0$ and $\gamma=0$ resembles the completely depolarizing channel used in Refs. [23, 18], which outputs the maximally mixed state in the infinite-dimensional Hilbert space, $\propto\sum_{k=0}^{\infty}|k\rangle\langle k|$, that is mathematically ill-defined and thus unphysical; however, because $\sigma$ is strictly larger than zero throughout the proof, the channels are well-defined. We will show that, even with the prior knowledge that the channel is from this family, it is still hard for entanglement-free schemes to achieve the learning tasks.
The key idea is to reduce learning to hypothesis testing. Consider the following game between Alice and Bob: Alice chooses one of two hypotheses with equal probability: (1) Set $\Lambda=\Lambda_{0}$; (2) Set $\Lambda=\Lambda_{\gamma}$ for $\gamma$ sampled from a Gaussian distribution whose variance is determined by $\kappa$. Next, Alice allows Bob to use the channel $\Lambda$ $N$ times, and Bob uses any entanglement-free schemes to learn from these channel uses. After Bob has finished all quantum measurements and keeps only classical data, Alice reveals some auxiliary information to Bob, and Bob then decides whether Alice has chosen (1) or (2).
Using a learning scheme satisfying the assumptions of Theorem 2, Bob can guess correctly with high probability; Thus, the outcome distributions of Bob’s scheme under hypotheses (1) and (2) must have a large total variation distance (TVD). Meanwhile, we prove that the contribution from each use of $\Lambda$ to the TVD is exponentially small using a new technique inspired by Ref. [45], which derived the maximum fidelity of Gaussian random displacement channels. Therefore, the number of channel uses $N$ must be exponentially large to ensure a large TVD, which gives us the desired lower bound on the sample complexity. ∎
@@
Effect of noise.— Now, for practical applications, we analyze the effect of dominant noise sources in optical platforms on the entanglement-assisted scheme and demonstrate its robustness against them, implying that the advantage is realizable shortly. We discuss the photon-loss effect in the main text, while phase diffusion and measurement crosstalk are covered in SM S2 D and E. Photon loss transforms the relevant bosonic operator $\hat{a}$ to $\sqrt{T}\hat{a}+\sqrt{1-T}\hat{e}$, where $T$ is the transmission rate and $\hat{e}$ is the environmental mode, i.e., $1-T$ is the loss rate. We consider two different places where the loss occurs: one is before applying the channel with loss rate $1-T_{b}$ to model the preparation imperfection, and the other is after applying the channel and before the perfect BM with loss rate $1-T_{a}$, which models an imperfect BM [27]. As before, we derive the relation between the measurement probability distribution and the characteristic function of the channel (with appropriate rescaling of the phase):
@@
which incorporates the loss rates. Because loss degrades the advantage from squeezing, the upper bound on sample complexity in Theorem 1 is modified as (see SM S2 C for the proof):